2. Aerobic degradation of an organic compound by a mixed culture of organisms in waste water can be represented by the following reaction. a. b. c. Determine a, b, c, d, and e, if Yxs-0.4 g X/gS. Determine the yield coefficients Yxio2 and YxNH3 Determine the degree of reductions for the substrate, bacteria, and RQ for the organism.
Answer #1

C=3.1=5.c+1.e

H=6.1+3.b=7.c+2.d

O=3.1+2.a=2.c+1.d+2.e

N=1.b=1.c

a)YX/S=o.4g-X/g-s=C.(5.12+7.1+1.14+2.16)/3.12+6.1+3.16

YX/S =0.4=C(113g/mol )/90=36=113.c

c=0.319

N: b=0.319

C: e=3-5.0.3/9=1.41

H: d=6+(3.0319)-7.0.319/z= 2.36

O: 3+2a=2.0.319+2.36+2.1.41

a=1.41

b=0.319

c=0.319

d=2.36

e= 1.41

   a=1.41

b)YX/O2 =0.319.113/1.41.325mol =YX/O2=0.80g/g

YX/NH3=0.319.113/0.319.17g/mol=6.65g/g=YX/NH3.

c)rs^1=3(4)+6(1)+(3-2) =12 rs=12/3=4

rb^1=5(4)+7(1)+2(-2)+1(-3)=20rb=20/5=4

RQ=mol CO2/molO2=e/a=1.41/1.41=1.0

rs=4

rb=4

RQ=1.0